Carla Vista Monograms sells stadium blankets that have been monogrammed with high school and university emblems. The blankets retail for $43 throughout the country to loyal alumni of over 3,800 schools. Carla Vista’s variable costs are 41% of sales; fixed costs are $118,000 per month.

a1) Calculate contribution margin ratio. (Round ratio to 2 percentage places, e.g. 0.38 = 38%.)

59%

a2) What is Carla Vista’s annual breakeven point in sales dollars? (Use the rounded contribution margin ratio calcuated in the previous part to compute breakeven sales.)

2,400,000

**b) Carla Vista currently sells 126,000 blankets per year. If sales volume were to increase by 15%, by how much would operating income increase? (Round answer to 0 decimal places, e.g. 5,275.)
Answer #1

a1.Variable cost per unit=(43*41%)=$17.63

Contribution margin=Sales-Variable cost  

=(43-17.63)=$ 25.37

Contribution margin ratio=Contribution margin/Sales

=25.37/43

=59%

a2.Annual fixed expenses= (118,000*12)=$1,416,000

Breakeven=Fixed expenses/Contribution margin ratio

=1,416,000/0.59

=$2,400,000

b.

Current Proposed
Blankets sold 126,000 (126,000*1.15)=144900
Contribution margin (126,000*25.37)=3,196,620 (144900*25.37)=3,676,113
Less:Fixed expenses 1,416,000 1,416,000
Operating income 1,780,620 2,260,113

Hence increase in operating income=2,260,113-1,780,620

=$479493